OpenStudy (anonymous):

WILL FAN AND MEDAL What is the height of a triangle with area 26 cm2 and base 8 cm?

OpenStudy (john_es):

The are of a triangle is, \[ A=\frac{base\cdot height}{2}\] Solve this equation for height and you will get the solution.

OpenStudy (anonymous):

hb=6.5cm Using the formula A=hbb 2 Solving for hb hb=2A b=2·26 8=6.5cm 6.5m